# The future of offshore wind power production: wake and climate impacts

### v1

@@ 4 Conclusions
Applying the Weibull quantile mapping bias correction method to downscaled future climate projections for the period 2070–2100, we then investigated and compared the impacts of climate change and wake effects, the latter evolving as a function of wind farm build-out in the region since 2010. Looking firstly at the climate effect, all climate models project decreases in wind power for the final 21-farm scenario for the 2070–2100 period compared with 1976–2006, with the magnitude of the change varying between 0.03 and 4.2%. Looking instead at wake effects and averaging over all climate models for the 2070–2100 period, we found that intra-cluster and inter-cluster wake effects grow from zero (by definition) when only one wind farm is present, to 4.3 and 2.5% respectively for the full 21-farm build-out scenario. Intra-farm wake-induced losses are less sensitive to the build-out, but are larger in magnitude, at 13.2% for the final build-out. Wake effects between farms and clusters have typically been underestimated in the context of resource assessment, but our results indicate that their impact after less than two decades of build-out is of greater magnitude than a century of climate change under the high-emissions RCP8.5 scenario. These wake effects are likely to continue to grow with further build-out, and it is vital to include their effects in long-term wind farm planning.
The potential to mitigate wake-induced losses in a robust way with respect to climate change requires further research. We note that there is limited agreement among the climate models in terms of which precise wind speeds and directions change in frequency in future. This lack of fine-scale agreement may prevent the design of specific farm layouts which are robust to future changes in wind climate, although it is likely that larger-scale design choices such as turbine sizes and densities can be informed by such projections. For example, reduced resource density in future may favour farms with greater turbine separation. Exploration of these aspects is left to future work.
We also observe evidence that the magnitude of wake effects differs between the historical and future climates, i.e. that climate and wake effects interact. Looking across all seasons, all climate models agree that there is a reduction in mean wind speeds, or equivalently an increase in frequency of lower speeds and a decrease in frequency of higher speeds. Since wake effects are strongest within a fairly narrow wind speed range, wake effects are projected to either increase or decrease depending on the precise wind speeds projected to increase or decrease in frequency. This high sensitivity leads to a lack of consensus between each climate model on the overall change in wake effects under the future climate scenario. However, isolating the summer months, all climate models agree that there will be an increase in the magnitude of wake-induced losses. This compounds the effects of seasonal variation (with lowest winds in summer), as well as the direct climate effect of weakening wind resources in the region. Given that wind farms are long-term investments, these results suggest that it is important to consider both climate change and wake effects at all spatial scales within wind farm planning.
This work has compared the overall magnitudes of climate and wake effects, but further work is needed to compare the uncertainties associated with each. This work has considered only a single climate scenario (which nevertheless carries significant uncertainty as demonstrated by the spread of results across the ensemble of downscaled climate projections). Uncertainty regarding future emissions will further increase the uncertainty in future wind power due to climate change. With regards future wake effects, in the short term it could be argued that these can be well constrained by data regarding planned wind farm developments, and thus carry little uncertainty. However, in the longer term, further uncertainty will also be introduced due to the unknown locations, sizes, performance characteristics and operational dates of future wind farms. The relative magnitudes of uncertainties introduced by climate change and future wind farm developments is not obvious and requires further research.
@@
In future work, we will also consider a broader variety of data sources for bias correction, such as other wind atlas products, and wind farm power output data. Such data sources may be valuable in overcoming the relatively coarse resolution of the ERA-5 reanalysis data used within this work, as well as helping to overcome any biases in this data. We will also consider investigating climate change on different time horizons and for different emissions scenarios, and using more recent projections such as are available from CMIP6. We will also look to include further planned wind farms over larger spatial regions. This will enable the assessment of the impact of the climate and wake effects on levelised cost of energy, and energy systems models.

### v2

@@ 5 Conclusions
Applying the Weibull quantile mapping bias correction method to downscaled future climate projections for the period 2070–2100, we then investigated and compared the impacts of climate change and wake effects, the latter evolving as a function of wind farm build-out in the region between 2010 and 2027. Looking firstly at the climate effect, all climate models project decreases in wind power for the final 21-farm scenario for the 2070–2100 period compared with 1976–2006, with the magnitude of the change varying between 0.03 and 4.2%. Looking instead at wake effects and averaging over all climate models for the 2070–2100 period, we found that intra-cluster and inter-cluster wake effects grow from zero (by definition) when only one wind farm is present, to 4.3 and 2.5% respectively for the full 21-farm build-out scenario, representing around two decades of build-out. Intra-farm wake-induced losses are less sensitive to the build-out, but are larger in magnitude, at 13.2% for the final build-out. Wake effects between farms and clusters have typically been underestimated in the context of resource assessment, but our results indicate that their impact after less than two decades of build-out is of greater magnitude than a century of climate change under the high-emissions RCP8.5 scenario. These wake effects are likely to continue to grow with further build-out, and it is vital to include their effects in long-term wind farm planning.
Our results indicate that wake-induced losses differ between the historical and future periods, under the RCP8.5 climate scenario. This implies an interaction between the climate and wake impacts. In spring and summer, wake-induced losses are found to be of greater magnitude under the future scenario, with the greatest change occurring during summer. This effect compounds existing seasonal variability in wind power, and exacerbates the negative impact of climate change on wind power within the German Bight region.
@@
In future work, we will consider a broader variety of data sources for bias correction, such as other wind atlas products, and wind farm power output data. Such data sources may be valuable in overcoming the relatively coarse resolution of the ERA-5 reanalysis data used within this work, as well as helping to overcome any biases in this data. We will also look to include further planned wind farms over larger spatial regions, and consider other regions worldwide where the relative magnitudes of climate and time-evolving wake effects may differ. This will enable the assessment of the impact of climate and wake effects on levelised cost of energy, and energy systems models.
@@ Acknowledgements
Appendix A Sensitivity to peak thrust coefficient
As described in section 2.1, the turbine power and thrust curves within this work were generated using PyWake’s ‘generic_wind_turbines’ module. While the power curve is typically well constrained by the turbine’s physical parameters (rotor area, cut-in and cut-out speeds, etc), the thrust coefficient ($C_{T}$) curve can exhibit greater variation between turbine models. A key parameter in constructing ‘generic’ $C_{T}$ curves is the peak $C_{T}$ value, which is assumed to be 0.8 by default. While this value is in reasonable agreement with publicly available turbine $C_{T}$ curves (NREL Turbine Archive, 2023), in this section we investigate the influence of this parameter on the key results from this study. To do this, we repeated the experiments presented in section 3.4 with peak $C_{T}$ values of 0.7 and 0.9. The results are summarised in table 1.
All wake effects increase in magnitude as the peak $C_{T}$ value increases. This is because the intensity and extent of turbine wakes increases with the turbine thrust. Larger peak $C_{T}$ values therefore induce greater wake losses. However, the overall trends are unchanged, and overall wake effects are still found to be of significantly greater magnitude than the mean climate impact, which has a value of -2.1%. Nevertheless, these results demonstrate the significance of uncertainty due to unknown turbine data. Studies of wake losses would benefit from greater sharing of such data by turbine manufacturers. However, these results also highlight the uncertainty inherent in modelling future wind farm scenarios, where even the basic turbine parameters may be unknown.
Peak $C_{T}$ value 0.7 0.8 0.9
Intra-farm wakes -12.2% -13.3% -13.8%
Intra-cluster wakes -4.0% -4.3% -4.4%
Inter-cluster wakes -2.3% -2.5% -2.6%
Total wake effect -18.4% -20.1% -20.8%
Table 1: Climate and wake effects for different values of the peak thrust coefficient ($C_{T}$), for the final build-out stage.
Appendix B Alternative climate scenarios
In this section we present results from analysing the RCP2.6 scenario, in contrast with the RCP8.5 scenario which is the focus of the main text. We also present results from analysing an intermediate future period of 2040–2070, compared with the period of 2070–2100 which is assessed in the main text.
While the RCP8.5 scenario is a high-emissions scenario sometimes referred to as ‘business as usual’, RCP2.6 is the lowest-emission scenario considered within CMIP5 (and hence EURO-CORDEX). Of the seven EURO-CORDEX models utilised to study the RCP8.5 scenario, data for the RCP2.6 scenario is available from only three. These are CNRM-A, NCC-R and MPI-R. For these three models, we repeated the experiments of section 3.3.2 using RCP2.6 data. For each model and climate scenario, table 1 summarises the projected change in mean power for the 2027 build-out stage, between the future (2070–2100) and historical (1976–2006) periods.
For the CNRM-A model, the projected impact of climate change on wind power is significantly greater for the RCP8.5 scenario than for the RCP2.5 scenario. However, for the NCC-R and MPI-R models, the impact is greater under the RCP2.6 scenario than RCP8.5. Averaging across the three models, there is a small increase in the climate impact between RCP2.6 and RCP8.5.
Increased impact from lower-emission scenarios is not necessarily surprising, given the complexity of the earth system and its (modelled) response to human activity. In particular, we note also that this study focuses on only a single location, hence the results will be sensitive to climate-induced changes in spatial patterns. This is also consistent with previous studies. For example, Fernández-Alvarez et al. (2023) downscaled projections from the CESM2 model, with their results showing differing signs of change in wind power density over the North Sea, depending on the climate scenario.
RCP2.6 RCP8.5
CNRM-A -2.3% -4.2%
NCC-R -1.1% -0.5%
MPI-R -1.0% -0.1%
Mean -1.4 -1.6%
Table 1: Comparison of climate change effects for the RCP2.6 and RCP8.5 scenarios. Only models for which both climate scenarios are available are shown.
Table 2 compares projected changes in wind power output for the 2027 build-out, for the period 2040–2070, and the period 2070–2100. We again find that there is a lack of agreement between climate projections as to which future period experiences the greatest decrease in wind power compared with the historical period. The mean effects are similar, although the 2070–2100 period is projected to produce slightly lower wind power than the 2040–2070 period. These results are again not unexpected; for example, Martinez et al. (2023) assessed changes in wind power density for Northern Europe including the North Sea, based on CMIP6 projections. Their results showed that the effect of climate change on wind power density is not necessarily monotonic with time.
2040–2070 2070–2100
CNRM-A -2.9% -4.2%
NCC-A -2.1% -2.2%
MOHC-A -2.5% -3.8%
CNRM-R -0.3% -0.9%
NCC-R -1.6% -0.5%
ICHEC-R -2.0% -2.9%
MPI-R -2.2% -0.1%
Mean -1.9% -2.1%
Table 2: Comparison of climate change effects for the RCP8.5 scenario, for the periods 2040–2070, and 2070–2100. For both time periods, the changes are calculated relative to the historical period.
Overall, the results of this section highlight the inherent uncertainty in climate projections, and the need for model ensembles in quantifying this uncertainty.